# Efficient Preimage Approximation for Neural Network Certification

### v1

@@ 3 Methods
The aim is to apply PREMAP [38, 39] on larger models with larger input spaces. This means we need to revisit some of the design decisions that worked well on smaller problems. In Section 3.1 we calculate tighter bounds for intermediate layers, in Section 3.2 we describe how we adaptively draw more Monte Carlo samples, and in Section 3.3 we design new heuristics for the selection of which neuron to split next.
@@ 3.1 Bound tightening
When a neuron is split, the linear relaxation error from that neuron towards all the following layers is reduced. However, $\beta$-CROWN [29], on which PREMAP is built, typically only updates the final layer because the optimization is computationally intensive [29]. Instead of using the full $\beta$-CROWN optimization, in this work, we propagate the bounds using backwards mode LiRPA [34]. These bounds are not as tight, but much faster to compute.
While LiRPA [34] only propagates bounds to the following layers in the network, a split also implies restrictions in previous layers. Hence, we derive how the bounds of previous layers and the input might tighten after a split. As with backwards mode LiRPA, these bounds are fast to compute but result in looser bounds.
@@
$[\underline{\mathbf{A}}^{lm}\mathbf{z}^{m}+\underline{\mathbf{b}}^{lm}]_{i}<0\textrm{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }0\leq[\overline{\mathbf{A}}^{lm}\mathbf{z}^{m}+\overline{\mathbf{b}}^{lm}]_{i}.$ (3)
@@
$\underline{\mathbf{z}}^{m}_{j}\leftarrow\max(\underline{\mathbf{z}}^{m}_{j},-(c+\underline{\mathbf{b}}^{lm}_{i})/\underline{\mathbf{A}}^{lm}_{ij})\leavevmode\nobreak\ \mid\leavevmode\nobreak\ \underline{\mathbf{A}}^{lm}_{ij}<0$
$\overline{\mathbf{z}}^{m}_{j}\leftarrow\min(\overline{\mathbf{z}}^{m}_{j},-(c+\underline{\mathbf{b}}^{lm}_{i})/\underline{\mathbf{A}}^{lm}_{ij})\leavevmode\nobreak\ \mid\leavevmode\nobreak\ \underline{\mathbf{A}}^{lm}_{ij}>0$
@@
$\underline{\mathbf{z}}^{m}_{j}\leftarrow\max(\underline{\mathbf{z}}^{m}_{j},-(c+\overline{\mathbf{b}}^{lm}_{i})/\overline{\mathbf{A}}^{lm}_{ij})\leavevmode\nobreak\ \mid\leavevmode\nobreak\ \overline{\mathbf{A}}^{lm}_{ij}>0$
$\overline{\mathbf{z}}^{m}_{j}\leftarrow\min(\overline{\mathbf{z}}^{m}_{j},-(c+\overline{\mathbf{b}}^{lm}_{i})/\overline{\mathbf{A}}^{lm}_{ij})\leavevmode\nobreak\ \mid\leavevmode\nobreak\ \overline{\mathbf{A}}^{lm}_{ij}<0$
for $\mathbf{z}^{l}_{i}\geq 0,$ where $c=\sum\nolimits_{k\neq j}\max(\overline{\mathbf{A}}^{lm}_{ik}\underline{\mathbf{z}}^{m}_{k},\overline{\mathbf{A}}^{lm}_{ik}\overline{\mathbf{z}}^{m}_{k})$. Note that the division introduces additional numerical instability, which we mitigate with padding around zero, making the bounds slightly looser. After the “reverse” propagation of bounds we use LiRPA to update the bounds of subsequent layers, as mentioned above. This happens in Algorithm 1 at the end of the SplitNeuron function on 9.
@@ 3.2 Sampling
The samples are used in PREMAP for three things: Monte Carlo estimation of the volume of both the preimage and the approximation (in Algorithm 2, 7 to 11), the gradient-based optimization of the bounds in $\beta$-CROWN to maximise the preimage approximation volume [39] (on 3 of Algorithm 2), and in the heuristics to select which neuron to split (in Algorithm 1, 8).
We could, like the previous version of PREMAP [39], only draw a large, fixed set of initial samples. Each split defines two subsets where we can reuse the samples (see 9 in Algorithm 1). However, the size of the subsets decrease exponentially with each split. Hence, in this work, when the subsets get too small we draw additional samples (7). At first, we use rejection sampling from uniform samples, $\mathbb{U}(\underline{\mathbf{x}},\overline{\mathbf{x}})$. However, just as with subsets, the hit rate decreases exponentially.
The inequalities introduced in Equation 3 create a polytope in the input space. To sample from this polytope we use hit-and-run sampling [21, 7]. The idea behind hit-and-run sampling is to start from a point inside the polytope, conveniently provided by 9 of Algorithm 1. Then we draw a random direction, calculate the distance to the edge of the polytope, and uniformly sample a new point on this line. Since consecutive steps are correlated, we take multiple steps between every recorded sample. This yields uniformly distributed samples [21] from the polytope and the procedure can be parallelized and GPU accelerated [7]. Note that Equation 3 is an outer bound, so we still need to do rejection sampling on the activations.
@@ 3.3 Heuristics
The heuristic for selecting which neuron to split is condensed in Algorithm 1 on 8. What we do is calculate a score for each neuron $i$ on layer $l$ and select the neuron with the highest score. The original PREMAP [39] aims to keep the splits balanced in terms of samples with a score function of
@@
where $\mathbf{z}_{ij}^{l}$ is the pre-activation for sample $j$ (out of $n$) at neuron $i$ on layer $l$. However, this heuristic offers no indication of how well LiRPA [34] might tighten the bounds. Therefore, we suggest replacing it with a combination of heuristics described below.
Since the layer $l$ is likely to be in the middle of the network, after a split the negative side, $\mathbf{z}^{l}_{j}<=0.0$, will be constant (by the definition of ReLU) while the positive side will be non-linear with respect to the output. Hence, as heuristics, we estimate how much the split would reduce the “uncertainty” on the negative part.
The heuristics are visualized in Figure 3. First, we calculate the linear bound with respect to the output, $f_{\mathbb{O}}(\mathbf{x})\geq\underline{\mathbf{A}}^{l}\mathbf{z}^{l}+\underline{\mathbf{b}}^{l}$ (or $f_{\mathbb{O}}(\mathbf{x})\leq\overline{\mathbf{A}}^{l}\mathbf{z}^{l}+\overline{\mathbf{b}}^{l}$ for an over-approximation). Then we calculate the area (yellow) that will be reduced,
$\sum_{k}\left|\underline{\mathbf{A}}^{lk}_{i}\underline{\mathbf{z}}^{l}_{i}\underline{\mathbf{z}}^{l}_{i}\right|/\max_{l^{\prime}i^{\prime}}\left(\sum_{k}|\underline{\mathbf{A}}^{l^{\prime}k}_{i^{\prime}}\underline{\mathbf{z}}^{l^{\prime}}_{i^{\prime}}\underline{\mathbf{z}}^{l^{\prime}}_{i^{\prime}}|\right),$
the maximum under-approximation (green),
$\sum_{k}\left|\underline{\mathbf{A}}^{lk}_{i}\underline{\mathbf{z}}^{l}_{i}\right|/\max_{l^{\prime}i^{\prime}}\left(\sum_{k}|\underline{\mathbf{A}}^{l^{\prime}k}_{i^{\prime}}\underline{\mathbf{z}}^{l^{\prime}}_{i^{\prime}}|\right),$
and the average extra uncertainty for the samples (blue),
$\sum_{k}\left(\sum_{j=1}^{n}\left|\underline{\mathbf{A}}^{lk}_{i}\min(\mathbf{z}^{l}_{ij},0)\right|/\sum_{j=1}^{n}\left[\mathbf{z}^{l}_{ij}<0\right]\right)/\max_{l^{\prime}i^{\prime}}(\ldots),$
where $k$ represents the, potentially, multiple outputs of $f_{\mathbb{O}}$. We combine these heuristics with a weighted sum that is empirically determined in Section 4.2. Before settling on these heuristics we also considered other options, which are documented in Appendix A.
$\mathbf{z}^{l}_{i}$$f_{\mathbb{O}}(\mathbf{x})^{k}$$\underline{\mathbf{z}}^{l}_{i}$$\overline{\mathbf{z}}^{l}_{i}$$\underline{\mathbf{A}}^{lk}_{i}\mathbf{z}^{l}_{i}$
Figure 3: Visualising the three new neuron selection heuristics. They are the yellow area, the green distance, and the average blue length.
It is possible that the selected split assigns all samples to one branch. This means either that (a) the other side is impossible to reach with the current constraints but the bounds given by LiRPA are not tight enough to show that, or (b) the other side is reachable, but our current estimate for that volume is zero. Hence, we can avoid processing that branch (in Algorithm 1, 10 or 11) without significantly affecting the volume of the under-approximation.
3.4 Parallelisation
In the original PREMAP [39, 38] generating the under- or over-approximation after a split (see Algorithm 1, 10 and 11) is done in parallel. Furthermore, $\beta$-CROWN [29] processes the branches (the while loop on 5) in batches to take advantage of GPU resources. For the pseudocode we opted not to show either for simplicity, but in practice we use both strategies.
Due to the early stopping criteria introduced by PREMAP (5 of Algorithm 1) batching has an additional advantage. If it is difficult to find tighter bounds for the primary branch selected by the heuristics (6 and 8), the “beam search” introduced by the batching might still progress and stop after fewer splits.
4 Experiments
In this section we empirically evaluate our improvements and demonstrate use-cases for PREMAP. In Section 4.1 we describe the experimental setup and datasets and in Section 4.2 we choose the weights for the new heuristics. We compare to the previous version of PREMAP in Section 4.3 and different network architectures in Section 4.4. In Section 4.5 we perform an ablation study and in Section 4.6 we use PREMAP as an investigation tool to enhance interpretability.
4.1 Experimental setup
The experiments were run with four cores from an Intel Xeon 6252 CPU and one NVIDIA RTX 2080 Ti, except for the control tasks in Section 4.3 that were run on an Apple M4 Pro CPU. All preimage approximations were limited to 10 minutes with a batch size of 2 (except the previous version of PREMAP that only processes the splits in parallel). Unless otherwise mentioned, we use a stopping volume threshold of $0.9$.
For most of the experiments we use the following image datasets: GTSRB [27] with traffic signs, SVHN [22] with house numbers, and CIFAR-10 [17] with general images. Details about the model architectures (fully connected and convolutional) and training is in Appendix B. For experiments with image datasets ($32\times 32$ pixels) we randomly sample images, evenly from every class, and patches with random sizes between $3\times 3$ and $6\times 6$ and random positions within the images. Additionally, in Section 4.3, we use five reinforcement learning control tasks further detailed in [39].
The code for the method and experiments will be released under an open source licence upon the acceptance of this paper.
4.2 Parameters
As described in Section 3.3 we now have three heuristics to combine. All of them are normalized to $[0,1]$ to make a weighted sum easier. For this experiment we sample images and patches as described in Section 4.1 and uniformly sample weights for the heuristics. The weights are sampled and evaluated multiple times so that we can normalize the times (by dividing by the mean). We then fit a Gaussian process with an RBF kernel and a white kernel and evaluate it on a grid. The top results can be seen in Table 1, where the “under” heuristic is slightly less important than “area” and “extra”.
Table 1: Using a Gaussian process to find weights for the heuristics. Showing the top results from a grid of $\{0.0,0.25,0.5,0.75,1.0\}$.
Time (s) area under extra
$0.715\pm 0.399$ $1.000$, $0.750$, $1.000$
$0.715\pm 0.400$ $1.000$, $1.000$, $1.000$
$0.718\pm 0.399$ $1.000$, $0.500$, $1.000$
$0.724\pm 0.399$ $1.000$, $0.250$, $1.000$
$0.726\pm 0.399$ $0.750$, $1.000$, $1.000$
4.3 Comparison
Here we compare our version of PREMAP with the previous version [39]. In Table 2 we consider the five smaller reinforcement learning tasks from [39], but run them with ReLU splitting instead of input splitting. In all but dubinsrejoin, the improved splitting heuristic leads to finishing within time and with fewer splits. The cartpole and dubinsrejoin preimages are harder to approximate without input splitting [39], but our new version makes more progress and processes more splits in the same time.
Table 2: Comparing our (new) version of PREMAP to the previous (old) version on five reinforcement learning controllers, averaging ten different seeds. The time limit is ten minutes with a stopping threshold of $0.9$ for the first two datasets and $0.75$ for the last three. We measure the coverage of the under-approximation (larger is better), the number of subdomains required to reach that (smaller is better), and the time it took (smaller is better).
Coverage Domains Time (s)
Dataset new old new old new old
auto_park $0.918$ $0.916$ $2.0$ $3.0$ $0.05$ $0.39$
vcas $0.931$ $0.905$ $17.0$ $21.1$ $0.16$ $5.57$
lunarlander $0.750$ $0.574$ $813.8$ $2745.8$ $36.60$ $600.15$
cartpole $0.750$ $0.424$ $1850.5$ $1601.7$ $118.05$ $600.14$
dubinsrejoin $0.207$ $0.000$ $4742.7$ $1159.0$ $596.11$ $600.16$
Table 3 contains results from the two versions of PREMAP on the GTSRB dataset. We aggregate the results from ten random images from every class with randomly placed and sized patches and count the number of preimage approximations that finish within the time limit. Some preimage approximations are easy enough to finish without entering the loop (Algorithm 1, 5 to 12), which is also visible in Figure 4. If we filter out those results, our version is able to complete more than twice the number of harder cases, especially for the convolutional neural network.
Table 3: Comparing our (new) version of PREMAP with the previous (old) version on the GTSRB dataset. Since some patch and image combinations reach the threshold of $0.9$ in a single iteration, the second row is filtered to exclude those. We measure the average coverage of the under-approximation (larger is better), the number of instances reaching the threshold within 10 minutes (larger is better), and the average time required (smaller is better).
Dataset: GTSRB Coverage Completed Time (s)
Model Subset # new old new old new old
FC All 500 $0.850$ $0.714$ $408$ $351$ $120.55$ $188.37$
FC Hard 292 $0.718$ $0.318$ $174$ $61$ $224.76$ $441.06$
CNN All 500 $0.751$ $0.501$ $358$ $244$ $186.91$ $313.70$
CNN Hard 342 $0.626$ $0.197$ $185$ $54$ $282.32$ $497.88$
4.4 Networks
In Figure 4 we run PREMAP on the image datasets using both fully connected (FC) and convolutional (CNN) neural networks. We divide the results into two groups based on how close to the edge the patch was. Both GTSRB and SVHN show strong dependence on the location of the patch, which is to be expected since the traffic signs and digits are in the centre of the images. Some SVHN images have distracting digits in the surroundings [22] making the robustness at the edges unsurprising.
Figure 4: Time to calculate a preimage under-approximation with a 10 minute time limit and a $0.9$ threshold. The patches are applied to random images (from every class) with random positions and sizes. We group the results based on if the centre of the patch is within 10 pixels of the centre of the image.
[…]
4.5 Ablation
In this section we perform an ablation study on a) the tightening of intermediate bounds and b) the new split selection heuristics. The result can be seen in Figure 5, where we measure the coverage of the under-approximation over time running on random patches, as before, and average the results. Tightening is better than using neither improvement, the heuristics are even better, and using both provides the best results. Note that the tightening of intermediate bounds introduces an additional computational overhead, which reduces the utility on a time versus coverage graph.
Figure 5: Comparing the average approximation coverage (higher is better) over time. The stopping threshold is $0.9$.
Figure 5 also demonstrates the anytime property of the algorithm: if we stop early we still get an under-approximation, just a smaller one, and when given more time more approximations reach the $0.9$ threshold.
4.6 Interpretability
A popular approach to explainable artificial intelligence (XAI) is to cover part of the image to see if the classification changes [14]. This reasoning is easy to understand, since if the prediction changes that area must have been important. However, this requires careful design to make a “neutral” replacement pattern [11], preferably such that the modified image is not be out-of-distribution [2]. With PREMAP we do not consider any singular replacement pattern, but rather the whole preimage of an area.
In Figure 6 we apply $4\times 4$ patches on a grid covering the image on the left. In the second image from the left we show the estimate of the volume of the preimage (not the approximation) for every patch, which indicates that this model is robust against patches of this size on this image. However, in the third image we see that patches around the ‘5’ require more splits for the approximation. This means that in these areas the model is highly non-linear. Furthermore, in the rightmost image we see that the preimage approximation of some patches do not reach a high coverage in the allotted time, notably around the areas where the ‘5’ could be turned into an ‘8’ with two patches. This is how PREMAP can be used to investigate models, which can aid model development and debugging [1].
Figure 6: Using PREMAP to investigate which parts of the image are important for classification without the need to provide a user-defined replacement pattern by applying patches on a grid covering the image. The figures are, from left to right, the original image, the estimated preimage volume (higher means more robust), the number of subdomains required (lower means easier), and the coverage of the preimage under approximations (higher is better).
5 Conclusions
This work introduced algorithmic improvements to the PREMAP algorithm for finding preimage approximations for neural networks. The main enhancements come from the improved heuristics for selecting splits, the calculation of tighter bounds after a split, and the adaptive Monte Carlo sampling. The improved performance is also empirically validated on both smaller control benchmarks and patch attack certification on images. We also demonstrate how to control the trade-off between time and approximation coverage and how PREMAP could be used for explainable AI. While we have focused our experiments on preimage under-approximation because of its inherent advantages for robustness and reliability certification, the methods also extend to over-approximations.
Future work includes applying PREMAP to other domains where certification is needed, for example biometric security or medical diagnosis. PREMAP could also be further enhanced to handle even larger and more complex networks, such as transformers. Finally, extending PREMAP to non-uniform input spaces would unlock additional applications, for example being able to prioritize dangerous conditions.
{ack}
This project received funding from the ERC under the European Union’s Horizon 2020 research and innovation programme (FUN2MODEL, grant agreement No. 834115) and ELSA: European Lighthouse on Secure and Safe AI project (grant agreement No. 101070617 under UK guarantee).
Appendix A Heuristics
To improve the heuristic for selecting which neuron to split we consider 12 alternative formulas for calculating the priority. These are detailed in Table 4. Note that not all are new, e.g., "balance" is the one used in the previous version of PREMAP [39, 38] and "lower" [3] and “gap” [23] are used in $\beta$-CROWN [29].
Table 4: Potential heuristics for selecting which neuron $i$ (on layer $l$) to split. There are $j\in[1,\ldots,n]$ samples and $k\in[1,\ldots,K]$ outputs. All formulas not already limited to $[0,1]$ are normalised by dividing by the maximum value and stable neurons with bounds $\underline{\mathbf{z}}^{l}_{i}\geq 0\vee\overline{\mathbf{z}}^{l}_{j}\leq 0$ are skipped.
@@ Appendix A Heuristics
lower $\max(-\underline{\mathbf{z}}^{l}_{i},0)$
@@
loose $\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i}-\bigl{(}\max_{j}(\mathbf{z}^{l}_{ij})-\min_{j}(\mathbf{z}^{l}_{ij})\bigr{)}$
bound $1-\bigl{(}\max_{j}(\mathbf{z}^{l}_{ij})-\min_{j}(\mathbf{z}^{l}_{ij})\bigr{)}/(\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i})$
gap $(-\underline{\mathbf{z}}^{l}_{i}\cdot\overline{\mathbf{z}}^{l}_{i})/(\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i})$
@@
stable $\min_{j}(\mathbf{z}^{l}_{ij})\geq 0\vee\max_{j}(\mathbf{z}^{l}_{ij})\leq 0$
pure $2\cdot\max\Bigl{(}\left|{\sum_{j}\bigl{[}\mathbf{z}^{l}_{ij}\geq 0\wedge f(\mathbf{x}_{j})\in\mathbb{O}\bigr{]}}/{\sum_{j}\bigl{[}\mathbf{z}^{l}_{ij}\geq 0\bigr{]}}-\frac{1}{2}\right|,$
$\left|{\sum_{j}\bigl{[}\mathbf{z}^{l}_{ij}<0\wedge f(\mathbf{x}_{j})\in\mathbb{O}\bigr{]}}/{\sum_{j}\bigl{[}\mathbf{z}^{l}_{ij}<0\bigr{]}}-\frac{1}{2}\right|\Bigr{)}$
@@
Some heuristics in Table 4 use the (linear) bounds produced by LiRPA [34, 29]. To minimize computational overhead we store them for reuse after each optimization. Some heuristics use the activation values of the samples. This requires computation on the scale of a forward pass through the network, which is not significant for the overall runtime.
To evaluate the heuristics we draw random samples (i.e., random images with random patches as in Section 4) and run PREMAP multiple times with random combinations of the heuristics and normalize the results by dividing by the mean times for that sample. The weights for the different heuristics are drawn such that roughly half are zero and the rest are uniform.

### v2

@@ 4 Methods
We aim to broaden the applicability of PREMAP [15, 16] to a range of use cases, which necessitates improvements in scalability and efficiency, including an extension to convolutional networks. Therefore, we will revisit some design decisions that worked well on smaller problems. In Section 4.1 we calculate tighter bounds for intermediate layers, in Section 4.2 we describe how we adaptively draw more Monte Carlo samples, in Section 4.3 we design new heuristics to select which neuron to split next, and in Sections 4.4 and 4.5 we describe computational strategies for increased efficiency.
@@ 4.1 Bound tightening
When a neuron is split, the linear relaxation error from that neuron to all the following layers is reduced. However, $\beta$-CROWN [27], on which PREMAP is built, typically only updates the final layer because the optimization is computationally intensive [27]. In this work, instead of using full $\beta$-CROWN optimization, for intermediate layers we propagate the bounds using simpler LiRPA [18]. These bounds are not as tight, but much faster to compute.
While LiRPA [18] only propagates bounds to the following layers in the network, a split also implies restrictions in previous layers. Hence, we derive how the bounds of previous layers and the input might tighten after a split.
@@
[…]
@@
$\underline{\mathbf{z}}^{m}_{j}\leftarrow$ $\max(\underline{\mathbf{z}}^{m}_{j},-(c+\underline{\mathbf{b}}^{lm}_{i})/\underline{\mathbf{A}}^{lm}_{ij})$ $~\mid~\underline{\mathbf{A}}^{lm}_{ij}<0$
$\overline{\mathbf{z}}^{m}_{j}\leftarrow$ $\min(\overline{\mathbf{z}}^{m}_{j},-(c+\underline{\mathbf{b}}^{lm}_{i})/\underline{\mathbf{A}}^{lm}_{ij})$ $~\mid~\underline{\mathbf{A}}^{lm}_{ij}>0$
@@
$\underline{\mathbf{z}}^{m}_{j}\leftarrow$ $\max(\underline{\mathbf{z}}^{m}_{j},-(c+\overline{\mathbf{b}}^{lm}_{i})/\overline{\mathbf{A}}^{lm}_{ij})$ $~\mid~\overline{\mathbf{A}}^{lm}_{ij}>0$
$\overline{\mathbf{z}}^{m}_{j}\leftarrow$ $\min(\overline{\mathbf{z}}^{m}_{j},-(c+\overline{\mathbf{b}}^{lm}_{i})/\overline{\mathbf{A}}^{lm}_{ij})$ $~\mid~\overline{\mathbf{A}}^{lm}_{ij}<0$
for $\mathbf{z}^{l}_{i}\geq 0,$ where $c=\sum\nolimits_{k\neq j}\max(\overline{\mathbf{A}}^{lm}_{ik}\underline{\mathbf{z}}^{m}_{k},\overline{\mathbf{A}}^{lm}_{ik}\overline{\mathbf{z}}^{m}_{k})$. Note that the division introduces additional numerical instability, which we mitigate with padding around zero, making the bounds slightly looser. After the “reverse” propagation of bounds we use LiRPA to update the bounds of subsequent layers. This happens in Algorithm 1 on line 11.
@@ 4.2 Sampling
The samples are used in PREMAP in three ways: for Monte Carlo estimation of the volume of both the preimage and the approximation (on line Algorithm 2, 7 to 11), the gradient-based optimization of the bounds in $\beta$-CROWN to maximize the preimage approximation volume [16] (on line 3 of Algorithm 2), and in the heuristics to select which neuron to split (in Algorithm 1, line 8).
We note that the original PREMAP [16] only draws a large, fixed set of initial samples. Each split defines two subsets where we can reuse the samples (see line 10 in Algorithm 1). However, the size of the subsets decrease exponentially with each split. Therefore, when the subsets get too small, we draw additional samples (line 7). At first, we use rejection sampling, however, the hit rate also decreases exponentially.
The inequalities introduced in Equation 3 create a polytope in the input space. To sample from this polytope we use hit-and-run sampling [34, 35]. The idea behind hit-and-run sampling is to start from a point inside the polytope, conveniently provided by line 10 of Algorithm 1. Then we draw a random direction, calculate the distance to the edge of the polytope, and uniformly sample a new point on this line. Since consecutive steps are correlated, we take multiple steps between every recorded sample. This yields uniformly distributed samples [34] from the polytope and the procedure is GPU accelerated [35]. Note that Equation 3 is an outer bound, so we still need to do rejection sampling on the activations.
@@ 4.3 Heuristics
In this section we discuss how to design and select appropriate heuristics to improve computational performance of preimage approximation, both for under- and over-approximations. The heuristic for selecting which neuron to split is condensed in Algorithm 1 on line 8, where a score is calculated for each neuron and the neuron with the highest score is selected. The original PREMAP [16] aims to keep the splits balanced in terms of samples with a score function that counts the number of samples on each side of the split:
@@
where $\mathbf{z}_{ij}^{l}$ is the pre-activation for sample $j$ (out of $n$) at neuron $i$ on layer $l$. However, this heuristic offers no indication of how well LiRPA [18] might tighten the bounds.
To improve the selection of which neuron to split, we consider alternative heuristics to calculate priority, detailed in Table I. We include the balance heuristic from the original PREMAP [16, 15] and the lower [23] and gap [36] heuristics used in $\beta$-CROWN [27]. All heuristics are normalized by dividing by the maximum value, unless already in $[0,1]$.
TABLE I: Heuristics for selecting which neuron to split
@@
lower [23] $\max(-\underline{\mathbf{z}}^{l}_{i},0)$
@@
loose $\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i}-(\max_{j}(\mathbf{z}^{l}_{ij})-\min_{j}(\mathbf{z}^{l}_{ij}))$
bound $1-(\max_{j}(\mathbf{z}^{l}_{ij})-\min_{j}(\mathbf{z}^{l}_{ij}))/(\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i})$
gap [36] $(-\underline{\mathbf{z}}^{l}_{i}\cdot\overline{\mathbf{z}}^{l}_{i})/(\overline{\mathbf{z}}^{l}_{i}-\underline{\mathbf{z}}^{l}_{i})$
@@
All formulas are for a single neuron (layer $l$, index ii), with $j\in[1,\ldots,n]$ samples in $\mathbf{X}$ and $k\in[1,\ldots,K]$ outputs of $f_{\mathbb{O}}$.
In Table I the balance [16] and soft heuristics prioritize neurons that would evenly split the samples. The lower [23] and width heuristics measure the bounds given by LiRPA, while loose and bound compare these bounds with the actual minimum and maximum values achieved by the samples. The gap [36] heuristic measures the vertical distance between the linear upper bound and origin in Figure 2.
As visualized in Figure 3, after a ReLU activation negative values $\mathbf{z}^{l}_{j}<=0.0$ are truncated while the positive values are non-linear with respect to the output of $f_{\mathbb{O}}$. Hence, the area, under, and extra heuristics estimate how much a split would reduce the “output uncertainty” on the negative part.
$\mathbf{z}^{l}_{i}$$f_{\mathbb{O}}(\mathbf{z}^{l}_{i})^{k}$$\underline{\mathbf{z}}^{l}_{i}$$\overline{\mathbf{z}}^{l}_{i}$$\underline{\mathbf{A}}^{lk}_{i}\mathbf{z}^{l}_{i}$
Figure 3: Visualizing three neuron selection heuristics: area (yellow area), under (green distance), and extra (the average blue length at the samples).
@@
Some heuristics in Table I use the (linear) bounds produced by LiRPA [18, 27]. To minimize computational overhead we cache them after each optimization. Other heuristics use the activation values of the samples, which requires a forward pass through the network. Hence, no heuristic significantly extends the overall runtime. The selection of which heuristics to use and their weights is performed in Section 5.2.
4.4 Shortcuts
In some cases, a branch can be discarded or finalized before running the heavy $\beta$-CROWN process, for example, when the branch contains no preimage samples (when under-approximating) or only preimage samples (when over-approximating). This lets us avoid some heavy computation (in Algorithm 1, on line 12 or 13) on that branch. For the remaining branch we can perform an additional split before running $\beta$-CROWN.
Furthermore, if all samples are assigned to one branch of a split it means that either (a) the other branch is impossible to reach, but the current LiRPA bounds are not tight enough, or (b) the other side is reachable, but our current estimate for that volume is near zero. We can detect these splits with just a single forward pass through the network, which lets us split and stabilize all of them simultaneously on line 9 of Algorithm 1 (before invoking $\beta$-CROWN).
4.5 Batching
The original PREMAP [16, 15] generates the approximations of a split in parallel (see Algorithm 1, 12 and 13). In addition, $\beta$-CROWN[27] processes the branches (the loop on line 5) in batches to take advantage of GPU resources. The pseudocode shows neither for readability, but in practice we use both strategies.
Due to the early stopping criteria introduced by PREMAP [16, 15] (line 5 of Algorithm 1), batching has an additional advantage: If it is difficult to find tighter bounds for the primary branch selected by the heuristics (6 and 8), the secondary branches in the batch might still progress and reach the threshold earlier.
5 Empirical evaluation
In this section we empirically evaluate the quality of our improvements to PREMAP. In Section 5.1 we describe the experimental setup and datasets and in Section 5.2 we choose the weights for the new heuristics. We compare to the original PREMAP on patch attacks and reinforcement learning tasks in Section 5.3 and on different network architectures in Section 5.4. In Section 5.5 we perform an ablation study.
5.1 Experimental setup
We ran all our experiments (including the use cases in Section 6) with four cores from an Intel Xeon 6252 CPU and one NVIDIA RTX 2080 Ti. All preimage approximations were limited to 10 minutes with a batch size of 2 (except the original PREMAP that does not support batching). Unless otherwise mentioned, we perform under-approximations with a stopping threshold for the approximation ratio of $0.9$ ($1.1$ for over-approximations).
We certify robustness against patch attacks on the following image datasets: GTSRB [19] (traffic signs), SVHN [20] (house numbers), and CIFAR-10 [21] (general images). Details about the model architectures (a fully connected network with three $300$ wide layers and a convolutional network with three convolutional layers and one fully connected layer) and training are included in Appendix B. For experiments with image datasets ($32\times 32$ pixels) we randomly sample images, evenly from every class. The patches are rectangular, with sizes between $3\times 3$ and $6\times 6$, and are randomly placed within the images.
In Section 5.3, we use the same five [37, 38, 39] (low-dimensional) reinforcement learning control tasks and networks from [16]. We compute preimage approximations for select actions (classes), like fire main engine in lunarlander and push left in cartpole, for the same input domains as [16] (with $x_{v}\in[-2,0]$} for lunarlander, $x_{\omega}\in[-2,0]$} for cartpole, and $x_{v}\in[-0.1,0.1]$} for dubinsrejoin).
For the use cases in Section 6 we, additionally, use the same synthetic dataset and network as in [14] for outlier detection and create a new dataset from GTSRB [19] and CIFAR-10 [21] for evaluating explanations.
The code for both the method and the experiments is available under an open source licence [^1] Available at: https://github.com/Anton-Bjorklund/Premap2.
5.2 Hyperparameters
To select which heuristics to use we sample images and patches as described in Section 5.1 and run the reinforcement learning control tasks with different seeds. For each task we uniformly sample multiple weights for the heuristics so that half are inactive and at least one has a weight of one:
$\boldsymbol{w}^{\prime}$ $\sim\mathbb{U}[-1,1]$
$\boldsymbol{w}$ $=\max(\boldsymbol{0},\boldsymbol{w}^{\prime})/\max(\boldsymbol{w}^{\prime}),$
and perform both over- and under-approximations.
To evaluate the heuristics we calculate the difference in the approximation ratio from after the first iteration (Algorithm 1, line 3) to after the final iteration (or 5 minute time limit) and divide by the total time:
$\delta=\left|\frac{v_{\mathbf{P}}(1)}{v_{\mathbb{O}}}-\frac{v_{\mathbf{P}}(n)}{v_{\mathbb{O}}}\right|/t(n).$ (4)
Larger values of $\delta$ indicate that the approximation ratio improves more quickly. To compare across tasks and datasets we normalize the $\delta$ values by dividing by the maximum $\delta$ for each task. We also filter out tasks where the heuristic choice makes no difference, such as when PREMAP finishes after the first iteration.
The results for individual heuristics can be seen in Figure 4, with a linear model showing the general trend. Enabling the “area”, “under”, and “extra” heuristics are correlated with faster improvements to the approximation ratio, while using “balance” [16] and “lower” [23] instead of another heuristic leads to slower approximations.